OpenStudy (anonymous):

What is the area of the sector bound by the center of the circle and arc CD in the circle below? 9.42ft^2 19.54ft^2 34.89ft^2 88.31ft^2

OpenStudy (anonymous):

@freckles can you help me with this?

OpenStudy (saifoo.khan):

Where's the circle?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Area of sector:\[\frac{\theta}{360} \times \pi r^2\]

OpenStudy (anonymous):

so it would be: 45/360 * pi(15)^2

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

ok thanks @saifoo.khan